### v2

# Sensing and Steering Stereotypes: Extracting and Applying Gender Representation Vectors in LLMs

Abstract
Large language models (LLMs) are known to perpetuate stereotypes and exhibit biases. Various strategies have been proposed to mitigate these biases, but most work studies biases as a black-box problem without considering how concepts are represented within the model. We adapt techniques from representation engineering to study how the concept of “gender” is represented within LLMs. We introduce a new method that extracts concept representations via probability weighting without labeled data and efficiently selects a steering vector for measuring and manipulating the model’s representation. We develop a projection-based method that enables precise steering of model predictions and demonstrate its effectiveness in mitigating gender bias in LLMs. [^1] Our code is available at: https://github.com/hannahxchen/gender-bias-steering
Sensing and Steering Stereotypes:
Extracting and Applying Gender Representation Vectors in LLMs
Hannah Cyberey, Yangfeng Ji, David Evans Department of Computer Science University of Virginia Charlottesville, VA 22904 {yc4dx,yangfeng,evans}@virginia.edu
1 Introduction
Large language models (LLMs) are optimized for making generalizations about the world based on their training data. These systems risk amplifying biases and inequities present in their training data, potentially perpetuating harmful stereotypes and resulting in discriminatory outcomes. To address these concerns, various mitigation strategies have been proposed, including techniques based on prompt engineering (Ganguli et al., 2023; Kaneko et al., 2024), fine-tuning (Chintam et al., 2023; Ranaldi et al., 2024), modified decoding (Lu et al., 2021; Liu et al., 2021), and detection (Inan et al., 2023; Fan et al., 2024).
While much research has explored gender bias in LLMs through a black-box approach, less attention has been paid to how these biases arise from the model’s internal workings. Recent work on representation engineering provides insights into varied abstract features within the internal representations of LLMs (Zou et al., 2023), such as sentiment (Tigges et al., 2023), spatiotemporal information (Gurnee and Tegmark, 2024), and true/false statements (Marks and Tegmark, 2024). Several studies have also demonstrated promising results in effectively controlling model behaviors by modifying their feature representations (Turner et al., 2023; Rimsky et al., 2024; Arditi et al., 2024).
In this work, we leverage activation steering (also known as activation engineering), to study how the concept of gender is encoded in the internal representations of LLMs affects their predictions and how we can manipulate internal representations to mitigate biases at inference time.
Figure 1: Steering “gender” concept in Qwen-1.8B, evaluated on an example from Winogenerated fill-in-the-blank task. Baseline shows the original probabilities with no steering applied.
Contributions. We propose a novel method that extracts linear representations from LLMs for steering model predictions associated with a given concept (Section 3). While existing methods for computing steering vectors rely on labeled data, we compute them using probability weighting without explicit data annotations. In addition, we introduce metrics to efficiently select a steering vector without exhaustive searches as was required by most previous methods. We show that steering vectors produced by our method exhibit a higher correlation with gender bias in model outputs than the prevailing difference-in-means method (Section 3.4). We then present an approach for applying steering vectors to provide precise control over the internal representation (Section 4). We demonstrate the effectiveness of our steering vectors and method for applying them in reducing gender bias on the in-distribution task (Section 4.2) and its potential to generalize to other application tasks (Section 4.3). Finally, we explore the generalization of our method for controlling bias associated with other protected attributes (Section 5).
2 Background
This section provides background on gender bias and activation steering for LLMs.
2.1 Gender Bias
The concept of gender is contested and multifaceted, encompassing a person’s self-identity and expression, the perceptions held by others, and the social expectations imposed upon them (Devinney et al., 2022). We adopt Ackerman (2019)’s definition of conceptual gender—the gender expressed, inferred, and used by a model to classify a referent through explicit (e.g., pronouns) or implicit associations (e.g., stereotypes). While some gender notions are multi-dimensional, we consider a simple setting where gender may be encoded in a one-dimensional subspace. We assume this subspace captures both explicit and implicit aspects that shape the model’s understanding of “gender”, such as explicit gender definitional terms and implicit gender traits or behaviors. Our work is grounded in gender schema theory (Bem, 1981), which describes the cognitive process of “gendering”—dividing entities into masculine and feminine categories—and its subsequent impact on individuals’ behaviors. We define gender bias as the prediction difference arising from conceptual differences in model representations of femininity and masculinity. This bias may or may not lead to undesirable outcomes (e.g., negative stereotypes and discrimination) depending on the context.
2.2 Activation Steering
Activation steering is an inference-time intervention that steers model outputs by deliberately perturbing the model’s activations (Turner et al., 2023). These activations (or residual stream activations) refer to the intermediate outputs aggregated from the preceding layers (Elhage et al., 2021). Model activations may be modified by applying steering vectors, which can be computed by different methods (Tigges et al., 2023) including logistic regression, principal component analysis, and difference-in-means (Marks and Tegmark, 2024) which is currently the most widely used method.
Consider a decoder-only transformer model, trained with a set of token vocabulary $\mathcal{V}$. The model makes predictions by mapping each input $x=(x_{1},x_{2},...,x_{t}),x_{i}\in\mathcal{V}$, to an output probability distribution $y\in\mathbb{R}^{|\mathcal{V}|}$. Given two sets of prompts, difference-in-means computes a candidate vector for each layer $l\in L$ as the difference in activation means:
$\vec{u}^{(l)}=\frac{1}{|\mathcal{D}_{A}|}\sum_{x\in\mathcal{D}_{A}}\vec{h}_{x_{i}}^{(l)}-\frac{1}{|\mathcal{D}_{B}|}\sum_{x\in\mathcal{D}_{B}}\vec{h}_{x_{i}}^{(l)}$
where $\vec{h}_{x_{i}}^{(l)}$ denotes the activation of input $x$ at token position $i$ and model layer $l$. The prompts in $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ are usually constructed with inputs reflecting two contrasting concepts. The vector $\vec{u}^{(l)}$ captures the internal representation difference between concepts $A$ and $B$ that may elicit changes in model outputs. While some work considers the last $n$ tokens, we follow most studies by computing vectors with only the activations at the final position.
Based on the candidate vectors of a size $|L|$, previous work often performs a brute-force search across layers to select the one with the optimal intervention performance (Arditi et al., 2024). During inference, the steering vector can be applied using activation addition (Rimsky et al., 2024), which intervenes in the forward pass of an input as:
$\vec{h}_{x}^{(l)}=\vec{h}_{x}^{(l)}+c\vec{u}^{(l)}$ (1)
where $c$ is the steering coefficient, which can be either positive or negative. This intervention is usually applied at the same layer from which the vector is extracted and across all input token positions.
3 Finding a Steering Vector
Our goal is to derive a steering vector that captures how the concept of gender is encoded in a model’s representation and that allows us to manipulate the internal representation’s gender signal in a controlled way. In this section, we introduce a method for extracting candidate vectors (Section 3.1) and an efficient approach for selecting the steering vector (Section 3.2). Section 4 discusses how we apply that steering vector at inference time.
3.1 Extracting Candidate Vectors
Let $A$ and $B$ denote two contrasting concepts (e.g., femaleness and maleness) each of which can be identified by an associated set of tokens. We measure the extent of $A$ and $B$ presented in a model for an input prompt $x\in\mathcal{D}$ based on its prediction output. We define the disparity score between the two concepts for an input $x$ as:
$s_{x}=P_{x}(A)-P_{x}(B)$
where $P_{x}(A)$ is the probability of predicting concept $A$ in the last token position output of $x$, aggregated over tokens for $A$. The disparity score indicates how likely an input would trigger the model to predict one concept over another in the next token prediction.
Let $f$ denote a function that maps each prompt $x\in\mathcal{D}$ to a partition as follows:
$f(x)=\begin{cases}\mathcal{D}_{A}&\text{if }s_{x}>\delta\\ \mathcal{D}_{B}&\text{if }s_{x}<-\delta\\ \mathcal{D}_{o}&\text{otherwise}\quad\quad(|s_{x}|\leq\delta)\end{cases}$
where $\delta$ is a score threshold that determines which concept the input is more likely associated with. Partition $\mathcal{D}_{o}$ represents neutral prompts that do not strongly relate to either concept.
In contrast to difference-in-means, which computes the activation mean difference between $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$, we incorporate neutral prompts with probability weighting to filter out signals unrelated to the target concepts. This allows the vector to capture a better representation of $A$ and $B$.
Suppose the average activation of neutral inputs $\mathcal{D}_{o}$ is $\vec{\bar{h}}^{(l)}_{o}$. For each layer $l\in L$, a candidate vector is computed as the weighted mean activation difference with respect to the neutral representations:
$\vec{v}^{(l)}$ $=\hat{\bm{v}}_{A}^{(l)}-\hat{\bm{v}}_{B}^{(l)}$ (2)
$\textrm{where}\quad\vec{v}_{A}^{(l)}$ $=\frac{\sum_{x\in\mathcal{D}_{A}}s_{x}(\vec{h}^{(l)}_{x}-\vec{\bar{h}}^{(l)}_{o})}{\sum_{x\in\mathcal{D}_{A}}s_{x}}$ (3)
We denote $\vec{h}^{(l)}_{x}$ as the activation of input $x$ in the last token position at layer $l$. The original input activations are position vectors measured from the origin of the latent space. However, this origin may differ from where the actual neutral position lies. To resolve this, we first offset each input activation $\vec{h}^{(l)}_{x}$ by the average neutral activations $\vec{\bar{h}}^{(l)}_{o}$. We then compute the aggregated vector representations for each concept by weighting the adjusted input activations by their corresponding disparity scores. The resulting candidate vector, $\vec{v}^{(l)}$, is simply the unit vector difference between $A$ and $B$.
3.2 Selecting a Steering Vector
We assume that the ideal vector would reflect the desired concept signal in both its direction and magnitude. It should be able to distinguish the concept that is more relevant to an input and to what extent. Under this assumption, we can evaluate the vectors similarly to a linear classifier. We compute a score using the projection measured on the candidate vector to classify each input. Given a separate set of prompts, $\mathcal{D}^{\prime}$, drawn from the same distribution as $\mathcal{D}$. We assess the linear separability of each candidate vector $\vec{v}\in\{\vec{v}^{(l)}\}_{l\in L}$ by the root mean square error (RMSE) as:
$\mathrm{RMSE}_{\vec{v}}=\sqrt{\frac{1}{|\mathcal{D}^{\prime}|}\sum_{x\in\mathcal{D}^{\prime}}\mathbb{I}_{\mathbf{sign}}(\mathbf{comp}_{\bm{v}}x\not=s_{x})\,s_{x}^{2}}$
where $\mathbf{comp}_{\bm{v}}x$ is the scalar projection of latent state activations $\vec{h}^{(l)}_{x}$ on vector $\vec{v}$ given input $x$. The indicator function $\mathbb{I}_{\mathbf{sign}}(\cdot)$ returns 0 if the scalar projection and disparity score of an input have the same sign, and $1$ if they have different signs. A vector $\vec{v}$ perfectly differentiates the concepts in direction when $\mathrm{RMSE}_{\vec{v}}=0$.
To evaluate how well a candidate vector captures the desired property, we compute the Pearson correlation between the scalar projection $\mathbf{comp}_{\bm{v}}x$ and the disparity score $s_{x}$ for each $x\in\mathcal{D}^{\prime}$. We select the final steering vector at the layer with the lowest RMSE score, excluding the 5% of the layers that are closest to the output (Arditi et al., 2024).
Figure 2: Candidate vector performance across model layers. The left y-axis shows the Pearson correlation between disparity scores measured in the model outputs and projections computed on the candidate vector. The right y-axis evaluates the linear separability for distinguishing the concepts, measured by the root mean square error (RMSE).
3.3 Experimental Setup
We test whether our method can find a steering vector that represents the concept of gender encoded in a model and is more effective than the prevailing method, difference-in-means (MD), in capturing this concept. We assume that gender is represented linearly along the dimension of feminine–masculine concepts, where we consider femaleness as concept $A$ and maleness as $B$ in our setup.
Dataset. The gendered language dataset consists of sentences generated by ChatGPT with gender-coded lexicons (Soundararajan et al., 2023), including adjectives that reflect stereotypical traits or characteristics of a certain gender (Gaucher et al., 2011; Cryan et al., 2020). Each sentence is labeled with the gender described and whether it is consistent with or contradictory to the gender stereotypes. As most sentences contain gender-definitional terms, we replace them with their neutral terms for half of the dataset. These sentences can help test the sensitivity of vectors to more neutral inputs that may or may not encode gender information. We split the dataset into a training set for vector extraction and a validation set for evaluating the vectors.
Models. We conduct the experiments with several popular open-source chat models (Qwen-1.8B and 7B, Llama-2-13B) and instruction models (Llama-3.1-8B, Granite-3.1-8B, Ministral-8B, Mistral-Nemo-12B, and OLMo-2-7B). Appendix F provides information about the references and model cards.
Our prompts ask the model to respond with the gender indicated in the given sentence, followed by a sentence from the dataset. Since some models do not directly respond with a gender-related token, we add an output prefix to guide the model to produce more relevant outputs in the next token prediction. For each gender concept, we randomly sample 800 prompts that satisfy the requirements of Section 3.1 for extracting the candidate vectors. The number of neutral prompts varies by model, but we subsample them if the size is larger than either the set of gendered prompts. We set the default score threshold $\delta$ to 0.05, but compare results using different $\delta$ values in Section E.2. Appendix A provides more details, including the gender tokens used for computing the disparity scores.
(a) Mistral-Nemo-12B-Instruct
(b) Qwen-7B-Chat
Figure 3: Disparity scores $s_{x}$ and scalar projections of each input from the validation set. The first and third columns show the baseline measured before intervention. The second and fourth columns illustrate the change in disparity scores by overlaying the results after debiasing from the left figures. The projections (x-axis) of all datapoints are measured before intervention. We perform intervention at the layer where the vector has the lowest RMSE.
3.4 Results
We evaluate the quality of candidate vectors extracted using our proposed weighted mean difference method (WMD) with the prior difference-in-means (MD) approach. Figure 2 shows the candidate vector performance on the validation set across all model layers, measured by RMSE and the projection correlation. Across all eight models we tested, both methods show a higher correlation between the vector projections and disparity scores and a lower RMSE score as the layer number increases. This suggests that the gender representations are generalized in later model layers. This aligns with previous findings that high-level concepts tend to emerge in middle to later layers (Zou et al., 2023; Rimsky et al., 2024). Results for other models are provided in Section B.1.
The best candidate vectors identified by WMD show a strong correlation with the disparity scores in model outputs and a high linear separability between the concepts of femaleness and maleness. We find that WMD maintains a consistently higher correlation than MD across six of the models, while showing a similar correlation for the other two models. The two methods show the largest performance gap for Qwen-7B, where the projection correlation of WMD is around 0.28% higher than the optimal layer of MD (Table 1). While both methods can identify layers with a low $\mathrm{RMSE}\approx 0$, the scores for WMD remain consistently lower than MD at layers with the highest correlation.
Figure 3 (first and third columns) compares the disparity scores and scalar projections measured for each input prompt with the steering vector selected at the optimal layer. Ideally, the projections should align closely with the green dashed line in the figure, reflecting a positive correlation with the disparity scores measured in model outputs. Our proposed method WMD yields a better correlation with the disparity scores, where inputs with a higher disparity show a larger projection value, as measured by the selected steering vector. It also reflects the degree of disparities more equally in both female and male directions. While MD does capture the gender representations to some extent, it poorly reflects with inputs more associated with the maleness concept where $s_{x}<0$, as shown in 3(b) for Qwen-7B model. For some of these inputs, the projections on the steering vector indicate a higher degree of female signal. This imbalance in generalization may impact their steering performance, which we demonstrate in the next section.
Baseline MD WMD Modal Interval
Model Bias Layer $r$ Bias Layer $r$ Bias
Llama-2-13B 0.49 29 0.81 0.28 37 0.85 0.16 [--0.33, 0.18]
Llama-3.1-8B 0.65 26 0.84 0.60 25 0.98 0.32 [--0.23, 0.15]
Ministral-8B 0.50 30 0.95 0.05 27 0.95 0.07 [--0.10, 0.12]
Mistral-Nemo-12B 0.65 35 0.89 0.08 37 0.98 0.02 [--0.32, 0.00]
Qwen-1.8B 0.53 19 0.88 0.14 19 0.88 0.14 [--0.95, 0.99]
Qwen-7B 0.51 26 0.69 0.32 29 0.88 0.12 [--0.27, 0.22]
Granite-3.1-8B 0.63 37 0.96 0.27 37 0.97 0.24 [--0.05, 0.05]
OLMo-2-7B 0.63 29 0.88 0.47 27 0.90 0.37 [--0.44, 0.16]
Table 1: Debiasing performance and projection correlation $r$ of the selected steering vector evaluated on the validation set. The bias score is the root mean square (RMS) of disparity scores. We report the bias score for the baseline model with no intervention and after applying steering vectors computed by MD and WMD. The layer indicates the layer number (from zero) from which the steering vector is selected.
4 Applying Steering Vectors
Previous works mostly consider contexts in which the model only needs to be steered in a particular direction or assume that the target directions are known in advance. However, in contexts such as bias mitigation, we need to apply steering based on the type of input, which may be unknown at deployment. We describe our method for applying the steering vector and demonstrate its efficacy in mitigating bias.
4.1 Intervention Method
Since a model can exhibit varied degrees of bias to different inputs and at different generation steps, we cannot achieve precise control of model behaviors by simply applying activation addition with a uniform steering coefficient (Equation 1). To obtain more precise control, we perform interventions for each input $x$ as follows:
$\vec{h}_{x}^{\prime}=\vec{h}_{x}-\mathbf{proj}_{\bm{v}}x+\lambda\cdot\hat{\bm{v}}$ (4)
where $\lambda$ is the steering coefficient and $\hat{\bm{v}}$ is $\vec{v}$ in unit vector form. When $\lambda=0$, it performs bias mitigation by subtracting the vector projection $\mathbf{proj}_{\bm{v}}x$, which captures the extent of bias in input $x$ reflected on the steering vector $\vec{v}$. We can steer the model outputs to either concept with a coefficient value of $\lambda\not=0$. The model becomes more biased to $A$ when $\lambda>0$ and to $B$ when $\lambda<0$. We apply this operation across all token positions of $x$ but at only the layer from which $\vec{v}$ was extracted.
This formulation is similar to directional ablation proposed by Arditi et al. (2024), which also considers vector projections. However, they show that this approach, using steering vectors computed by MD, can only be used for removing a single concept (in one direction) and requires interventions across all model layers. Our proposed intervention provides a unified formulation for concept removal and steering model behaviors in either direction.
4.2 Steering for Bias Mitigation
We evaluate the effectiveness of steering vectors selected using the method described in Section 3.4 to mitigate gender bias. We apply the steering vectors with our proposed projection-based intervention with $\lambda=0$ and measure the bias score on the validation set, computed as the root mean square (RMS) of disparity score $s_{x}$.
Table 1 reports the bias scores before and after debiasing for each model. After applying the intervention, it shows a significant reduction in the bias score for all models. The intervention is particularly effective for Ministral-8B and Mistral-Nemo-12B instruction models with bias scores reduced to nearly zero. In addition, the results suggest that the projection and bias score correlation $r$ is a good indicator of the intervention performance. Models with a higher value of $r$ show a greater decrease in the bias score after intervention.
To analyze the impact of intervention on different inputs, we compare the bias score difference and the scalar projection of each input, as shown in the second and fourth columns of Figure 3. We apply the same intervention method for both steering vectors computed by MD and WMD. The projections of all data points are measured on the baseline model with no intervention. Debiasing with WMD’s steering vectors works as intended, where more biased inputs show a larger difference in their bias scores after debiasing, while less biased inputs are less affected. However, the inputs tend to be over- or under-corrected in their bias scores when using steering vectors computed by MD. As our intervention approach depends on the projection of each input, the mitigation becomes less effective when the steering vector fails to separate the bias direction or does not reflect well with model bias.
4.3 Steering Transferability
We evaluate the robustness of steering vectors computed using our method by testing whether a steering vector extracted using one dataset transfers effectively to other tasks.
4.3.1 Evaluation Tasks
We consider two gender bias tasks:
Winogenerated (Perez et al., 2023) is a human validated version of the Winogender pronoun resolution task (Rudinger et al., 2018) that is 50 times larger than the original datset. The model is asked to fill in the missing blank with a pronoun for a given sentence (e.g., “The surgeon assured the patient that __ would do the best possible job.”). The response can be either a male, female, or gender-neutral pronoun. We report the output probability produced for each gender pronoun, normalizing over all three pronoun options.
Occupational Stereotypes. We construct a question-answering style task that asks the model, What does [name] work as at the [industry/place]?. We use terms from nine different industries (e.g., technology, healthcare) and 100 first names commonly associated with each female, male, and gender-neutral group. We measure the frequency of job titles mentioned in the model’s generated response for each group under the model’s default temperature setting. Note that the prompts do not contain any explicit gendered words except for names that may encode gender information.
Appendix C provides further details on the construction of both tasks.
4.3.2 Results
Figure 4: Average output probabilities for male (M), female (F), and neutral (N) pronouns. The shaded areas show the standard deviation from the average. Results shown are based on steering Qwen-1.8B over 1.2K Winogenerated examples.
We test whether the same steering vector, extracted from the gendered language dataset, can be applied to manipulate gender signals in the model for different tasks. We apply the intervention approach described in Section 4.1 with different steering coefficients $\lambda$ on the Winogenerated task. Figure 1 shows an example of output probabilities produced by steering Qwen-1.8B. In Figure 4, we show the overall output probabilities based on the average of 1.2K randomly sampled examples from the dataset.
When $\lambda=0$, we remove bias by the degree of gender signal reflected on the steering vector (Equation 4). As shown in Figure 4, when $\lambda\simeq 0$, our method effectively reduces gender bias in the model, with neutral pronouns having the highest probability, while male and female pronouns show similar but lower probabilities on average. The effect of coefficient values on the model’s outputs also aligns with the expected gender concept. A more positive $\lambda$ increases the output probability for female pronouns, whereas a more negative $\lambda$ increases it for male pronouns. The model is less likely to predict neutral pronouns when steering with a larger magnitude of $\lambda$ in either direction.
For the occupational stereotypes task, we analyze the frequency difference in job titles predicted for feminine and masculine names before and after debiasing with steering vectors. Figure 5 displays the predicted job titles in the technology and healthcare sectors with the largest gender disparities. Prior to intervention, the model exhibits the largest discrepancies in predicting “software engineer” and “product manager” in technology and “nurse” and “doctor” in healthcare. Debiasing substantially decreases the frequency gap for most common job titles, and increases the relative prediction frequency of more neutral titles, such as “healthcare professional” for masculine names.
Figure 6 reports the distribution of scalar projections measured from prompts for five industries. Despite the lack of explicit gender wording in prompts, the projections measured indicate that the model still infers gender signals from the input. The projections also correspond to the gender associated with the names provided in the prompts. Masculine names show higher negative projection values, while feminine names exhibit higher positive projections. Gender-neutral names tend to have the lowest magnitude of projections. This shows the potential of using steering vectors to detect implicit gender bias in models that may be difficult to identify through black-box evaluation.
Figure 5: Difference in job title prediction frequency when prompted with feminine names compared to masculine names. The color represents the difference before and after debiasing on Qwen-1.8B. The y-axis shows the top 10 titles with the largest prediction gap.
Figure 6: Scalar projections of examples from the occupational stereotypes task, evaluated on Qwen-1.8B at the last token position. The color indicates the gender associated with the name used in the prompt.
5 Steering Racial Bias
This section explores generalizing our method to other protected attributes. We experiment with the concept of racial majority–minority groups, where the majority is White American and the minority is Black American. [^2] As with gender, race is a complex and non-binary notion that cannot be fully captured with a single dimension. We do not intend to suggest any kind of racial binary by using these categories, just select these as representative categories to enable our experiments because of the availability of data from previous linguistic experiments. We show that our proposed mitigation can be applied similarly to reduce racial bias in models.
5.1 Setup
We apply the approach introduced in Section 3 to find a steering vector for manipulating white and black racial concepts in the model. We use two dialectal datasets with written sentences in White Mainstream English (WME) and African American Language (AAL) [^3] We follow the terminology used by Lanehart et al. (2015) and provide more background in Section D.1.: (1) Groenwold et al. (2020) includes paired AAL texts from Twitter and WME equivalents translated by humans. (2) Mire et al. (2025) contains machine-translated AAL instructions from RewardBench (Lambert et al., 2024), which aligns more with WME. These datasets are different from the gendered language dataset that contains third-person descriptions with explicit gender markers (Section 3.3). We hypothesize that the steering vector can be captured by the sociolinguistic differences between WME and AAL speakers.
We construct prompts that ask for the most likely race based on the dialect of a sentence randomly sampled from the datasets. We compute the disparity score based on the model’s output probability of race-associated tokens (e.g., White, Caucasian, Black, African). A disparity score $s_{x}>0$ suggests the input $x$ is more associated with black, whereas $s_{x}<0$ indicates a higher white signal is presented in $x$. Section D.2 provides more details of the experimental setup.
5.2 Results
7(a) compares the disparity scores before and after mitigating racial bias with the steering vectors we found for Llama-3.1-8B and Mistral-Nemo-12B. The steering vectors for both models show a strong correlation with the disparity scores before debiasing. In 7(b), we compare the model’s output probabilities for both racial concepts when applied with different steering coefficients $\lambda$. The probabilities (as shown by the solid lines) are measured by the normalized output probabilities of white- and black-associated tokens, averaged over 200 sampled inputs. The result after debiasing in 7(a) corresponds to $\lambda=0$ in 7(b). Both models show a similar probability between white and black, which aligns with our intended goal of debiasing. The effect of the coefficient value $\lambda$ is also consistent with the desired model behavior. A higher positive value increases the probability of predicting black-associated tokens, whereas a larger negative $\lambda$ increases the probability of predicting white-associated tokens.
Our results demonstrate how our proposed method can be generalized to manipulate other protected attributes and mitigate bias in models. Additional results are provided in Section D.3.
(a) Disparity scores before and after debiasing.
(b) Racial concept probability with varied coefficients $\lambda$.
Figure 7: Steering racial concepts in Llama-3.1-8B and Mistral-Nemo-12B. All results are measured on the validation set. (a) All projections are computed before intervention. (b) The probability for each concept is averaged over 200 randomly sampled examples.
6 Conclusion
This paper introduces a new method for computing steering vectors to control model outputs related to a specific concept. We demonstrate its effectiveness in finding gender steering vectors that exhibit a stronger correlation with the gender concept compared to the widely-used method. Further, we present a technique for applying this steering vector to reduce gender bias in model prediction. Our results show that we can apply steering vectors extracted using our method to precisely decrease bias for the in-distribution task and that the extracted vectors are general enough to achieve promising results when transferred to different tasks. In addition, our method can be applied similarly to manipulate other types of protected features.
Limitations
Our work studies gender representations in LLMs, specifically through the feminine–masculine spectrum. We acknowledge the limited scope of our approach, as it examines gender through a single dimension, which oversimplifies the complex, multifaceted nature of gender identity and expression. Moreover, our emphasis on the binary spectrum fails to account for non-binary and fluid gender identities. Another critical limitation relates to the phenomenon of fairness gerrymandering (Kearns et al., 2018), which suggests models may appear to be fair along individual demographic dimensions while exhibiting biases against intersectional subgroups. Our one-dimensional approach may mask disparities affecting the intersection of multiple demographic dimensions. While our initial results on the transferability of steering vectors are promising, they require further rigorous testing. Future research should expand the scope of evaluation to a broader range of tasks and adopt a more comprehensive approach that considers the intersectionality of gender with other social identities.
Appendix A Prompt Construction Details
We split the gendered language dataset in half into training and validation sets. The results presented in Section 3.4 and Section 4.2 are based on 1600 prompts randomly sampled from the validation split.
A.1 Target Gender Tokens
We construct the set of target gender tokens based on the definition of “linguistic gender’ (Cao and Daumé III, 2020)’ and “generic pronouns” (Doughman et al., 2021) as follows:
  • Female: female, females, feminine, she, her, hers, woman, women, girl, girls, ♀
  • Male: male, males, masculine, he, his, him, man, men, boy, boys, ♂
We also include tokens of their uppercase and ones added with a prefix space (e.g., “ she”)
A.2 Prompt Instructions
Training set [Instruction | Output prefix]
What is the gender mentioned in the following? | The gender mentioned is
Determine the gender identity described in the following sentence. | The gender identity described in the sentence is
Identify the gender described in the following. | The gender described is
Output the gender suggested in the following. | The gender suggested is
Output the gender identity indicated in the sentence. | The gender identity indicated in the sentence is
Output the gender encoded in the text. | The gender encoded in the text is
Respond with the embedded gender in the text that follows. | The gender embedded in the text is
Reply with the gender denoted by the text as follows. | The gender denoted by the text is
Answer with the gender depicted in the sentence. | The gender depicted in the sentence is
Return the gender specified in the text. | The gender specified in the text is
Return the embedded gender in the text that follows. | The embedded gender in the text is
Provide the gender implied in the following text. | The gender implied in the text is
Provide the gender reflected in the following sentence. | The gender reflected in the sentence is
State the gender identity inferred from the sentence. | The gender identity inferred from the sentence is
State the gender implied in the text. | The gender implied in the text is
Specify the gender suggested in the text. | The gender suggested in the text is
Reveal the gender inferred in the text. | The gender inferred in the text is
From the following sentence, indicate the gender identity. | The gender identity in the sentence is
Based on the following details, reply with the gender implied. | The gender implied is
Based on the information provided, state the associated gender identity. | The gender identity in the provided information is
Validation set [Instruction | Output prefix]
Determine the gender entailed in the text. | The gender entailed in the text is
Determine the gender identity suggested in the sentence. | The gender identity suggested in the sentence is
Identify the gender indicated in the statement. | The gender indicated in the statement is
Output the gender suggested in the sentence. | The gender suggested in the sentence is
Output the gender inferred in the text. | The gender inferred in the text is
Respond with the gender specified in the text that follows. | The gender specified in the text is
Answer with the gender denoted below. | The gender denoted is
Return the gender portrayed in the sentence. | The gender portrayed in the sentence is
Provide the gender described in the following text. | The gender described in the text is
State the gender denoted in the text. | The gender denoted in the text is
Reply with the gender mentioned in the text. | The gender mentioned in the text is
From the following sentence, indicate the gender identity. | The gender identity described in the sentence is
Based on the following, respond with the associated gender. | The gender associated with the text is
Based on the given information, output the gender depicted. | The gender depicted in the given information is
Appendix B Steering Gender Bias
B.1 Candidate Vector Performance
Figure 8: Candidate vector performance across model layers. The left y-axis shows the Pearson correlation between disparity scores measured in the model outputs and projections computed on the candidate vector. The right y-axis evaluates the linear separability for distinguishing the concepts, measured by the root mean square error (RMSE).
B.2 Bias Mitigation with Steering Vectors
Figure 9: Disparity scores before and after debiasing the model with the steering vector. The x-axis indicates the scalar projection of each input before intervention.
B.3 Steering Coefficient vs. Gender Probability
Figure 10: Average output probabilities of female and male gender concepts with varied coefficients, evaluated on 200 randomly sampled examples from the validation set. The probability is measured based on the set of target tokens associated with either gender.
Appendix C Steering Transferability
C.1 Prompt Construction
[…]
Figure 11: Prompt template used for the Winogenerated task.
Instruction: Question: What does [NAME] work as at the [INDUSTRY/WORKPLACE]? Response: Answer: [NAME] works as a
Figure 12: Prompt template used for the Occupational Stereotypes task.
Term replacements for [INDUSTRY/WORKPLACE], categorized by industries:
  • Technology: tech industry, technology sector, information technology industry, IT industry
  • Legal: legal industry, law sector, legal service industry
  • Healthcare: hospital, healthcare industry, pharmaceutical industry
  • Public: government sector, public sector, state government, public services industry
  • Education: education industry, educational services sector, education sector
  • Sales: retail industry, marketing industry, sales industry, commercial industry
  • Finance: financial sector, finance industry, business sector, financial services industry
  • Entertainment: media industry, media sector, entertainment industry
  • Industrial: manufacturing industry, industrial sector, transportation industry
C.2 Winogenerated
Figure 13: Average output probabilities for male (M), female (F), and neutral (N) pronouns. The shaded areas show the standard deviation from the average. Results shown are based on steering Ministral-8B over 1.2K Winogenerated examples.
C.3 Occupational Stereotypes
Figure 14: Top predicted tokens of Qwen-1.8B with varying coefficients given an example from the occupational stereotypes task. The output probabilities are normalized over the tokens listed.
Figure 15: Difference in job title prediction frequency when prompted with feminine names compared to masculine names. The color represents the difference before and after debiasing on Qwen-1.8B-Chat. The y-axis shows the top 12 titles with the largest prediction gap.
Figure 16: Input projections of the occupational stereotypes task, evaluated on Qwen-1.8B-Chat at the last token position. The color indicates the gender associated with the name used in the prompt.
Appendix D Steering Racial Bias
D.1 English Dialects: WME vs AAL
White Mainstream English (WME) or Standard American English (SAE) is the dominant English dialect spoken by most people across the United States. We consider African American Language (AAL) as another English variety in our experiment, given its extensive literature in sociolinguistics. Other names have also been used in the literature, including African American English (AAE) and African American Vernacular English (AAVE). AAL has unique phonological and grammatical patterns that distinguish it from WME. Previous work has shown that NLP models tend to favor dominant languages like WME but exhibit poor performance when prompted with other English dialects, such as AAL (Fleisig et al., 2024).
D.2 Setup
Target Tokens. We use the following tokens for White and Black racial groups: White, Caucasian, Black, African (for African American). We include a few other variations of the tokens, using a similar approach described in Section A.1.
Prompt Templates [Prompt | Output prefix]:
  • What is the most likely race of a person who says “[INPUT]” based on the dialect? | The race is most likely to be
  • Based on the dialect, answer directly with the racial identity of a person who said “[INPUT]”. | The racial identity is likely to be
  • Based on the dialect, output the most likely race of a person who says “[INPUT]”. | The race is most likely to be
  • Based on the dialect, what is the most likely race of someone who says “[INPUT]”? | The race is most likely
We extract steering vectors with 1,000 WME and 1,000 AAL prompts and perform evaluation on a separate set of 1.2K randomly sampled prompts. We set the threshold score $\delta$ to 0.1 to identify prompts for each racial concept. We notice that the disparity scores for some models are heavily skewed towards either race. Therefore, we perform weighted sampling by the inverse square of the disparity score bin frequency ($1/n^{2}$) to sample prompts for extracting vectors.
D.3 Additional Results
(a) Disparity scores before and after debiasing.
(b) Racial concept probability with varied coefficients $\lambda$.
Figure 17: Steering racial concepts in Qwen-7B-Chat. We evaluate on the validation set with intervention at layer 26. (a) The projections shown are measured before intervention. (b) The average probability (solid line) is computed over 200 randomly sampled examples.
Appendix E Analysis
This section analyzes the impact of disparity score distribution and the choice of score threshold $\lambda$ on the resulting steering vectors’ quality and intervention performance.
E.1 Impact of Disparity Score Distribution
Figure 18: Probability distribution of disparity scores over the entire training set from which the prompts used for extracting vectors are sampled.
We analyze how the disparity scores of the training set for extracting vectors may impact the quality and intervention performance of steering vectors. Figure 18 shows the disparity score probability distribution over the entire training set for each model. Most models exhibit a similar tri-modal distribution pattern with three distinct peaks located around -1, 0, and 1, except for Qwen-1.8B, which shows a unimodal distribution. This demonstrates these models’ ability and tendency for “gendering” texts into female and male categories. We compute the mode intervals of the distribution using the SkinnyDip algorithm (Maurus and Plant, 2016), based on the dip test of unimodality (Hartigan and Hartigan, 1985), as shown by the shaded areas in Figure 18. Our results suggest that models with a wider center modal interval, like Llama-3.1-8B and OLMo-2-7B, show less effective debiasing performance with steering (Table 1). Furthermore, we find that models with less prominent peaks in their distribution, such as Llama-2-13B and Qwen, also show a lower projection correlation in their steering vectors.
E.2 Varying Disparity Score Threshold
Figure 19: Bias scores after intervention using steering vectors computed by eight different threshold scores for constructing the training set, where $\delta=[0.01,0.3]$.
Results shown in both Section 3.4 and Section 4.2 are based on the same score threshold $\delta$ of 0.05. We test the robustness of both vector extraction methods under different threshold values and measure their resulting steering vectors’ debiasing performance on the same validation set. We use eight different values of $\delta$ from 0.01 to 0.3 with increasing increments. Figure 19 shows the range of RMS bias scores after debiasing under different $\delta$ across all eight models. achieves comparable debiasing effects across all models, with a difference of less than 0.05 in bias scores for the same model. MD exhibits the largest discrepancy in bias scores for the Llama-3.1-8B model, with a difference of 0.1. While MD does not show a significant change in bias scores for most models, the bias scores consistently remain higher than those of WMD after debiasing.
Appendix F Models
Model Reference Model Card
Qwen-1.8B Bai et al. (2023) Qwen/Qwen-1_8B-Chat
Qwen-7B Qwen/Qwen-7B-Chat
Llama2-13B Touvron et al. (2023) meta-llama/Llama-2-13b-chat-hf
Llama3-8B Dubey et al. (2024) meta-llama/Llama-3.1-8B-Instruct
Ministral-8B Mistral AI team (2024b) mistralai/Ministral-8B-Instruct-2410
Mistral-Nemo-12B Mistral AI team (2024a) mistralai/Mistral-Nemo-Instruct-2407
Olmo2-7B OLMo et al. (2024) allenai/OLMo-2-1124-7B-Instruct
Granite3.1-8B Granite Team (2024) ibm-granite/granite-3.1-8b-instruct
Table 2: Model cards used in the experiments.
Appendix G Generated Outputs Before and After Steering
[…]
G.1 Steering Gender
Example G.1  (Baseline) Instruction: Imagine a woman whose name is Tiffany. Please describe her. Response: Tiffany is a young woman who has bright, sparkling eyes that seem to sparkle with happiness and joy whenever she’s around. She has long, curly brown hair that falls in loose waves around her shoulders, and she often wears natural, earthy makeup that accentuates her features without being too bold or heavy. Tiffany is tall and slender, with a lean build that gives her a dynamic, athletic appearance. She has a contagious energy about her that seems to fill the room around her, and she exudes kindness and generosity to everyone she meets. When she’s not working or studying, Tiffany enjoys spending time outdoors. She loves hiking, camping, and exploring new places. She also enjoys reading and learning, whether it’s through books, online courses, or attending workshops and seminars. Despite her busy schedule, Tiffany always manages to make time for her loved ones. She is a devoted partner and mother, and she spends countless hours nurturing her relationships with her husband, children, and siblings. She is also an active member of her community, volunteering at local organizations and participating in charity events. Overall, Tiffany is a bright, ambitious, and compassionate person who makes a positive impact on those around her. Her infectious enthusiasm and infectious laughter are sure to brighten up anyone who meets her.
Example G.1   Steering Output ($\lambda=-20$) Tiffany is a tall, athletic man with curly brown hair and piercing blue eyes. He has a lean build and is dressed in well-fitted clothing that emphasizes his strong physique. Despite his rugged exterior, he exudes a sense of confidence and strength. In terms of personality, Tiffany is outgoing and adventurous. He enjoys trying new things and exploring different cultures, which often leads him to meet interesting people and try new experiences. He is also highly intelligent and analytical, always seeking to understand the world around him. Outside of work, Tiffany is a devoted husband and father. He spends most of his time with his family, playing sports together, and cooking delicious meals. He is also an avid reader and enjoys staying up late discussing books and philosophy with friends. Overall, Tiffany is a dynamic and charismatic individual who brings energy and excitement to all those around him.
Example G.1   Steering Output ($\lambda=0$) Tiffany is an 25-year-old American professional athlete who competes in various sports such as basketball and football. She stands at 6 feet tall and has short, curly brown hair that falls in loose waves around her shoulders. In terms of physical appearance, Tiffany is muscular and athletic, with broad shoulders and a lean build. Her skin is smooth and tanned from years of exercise and training, and she always wears workout clothes that are comfortable and functional. When it comes to her personality, Tiffany is confident, competitive, and driven. She works hard to achieve her goals and takes pride in her achievements both on and off the field. Despite her success, she remains humble and down-to-earth, always making time for her friends and family and striving to be a positive role model for others. Overall, Tiffany is a dynamic and inspiring young woman who embodies the spirit of sports and competition.
Example G.1   Steering Output ($\lambda=30$) Tiffany is an elegant and graceful young woman with long, flowing hair that cascades down her back in soft waves. She has big brown eyes that sparkle with intelligent curiosity and radiate kindness towards those around her. Tiffany has a vibrant personality and a contagious energy that brings joy to the lives of those she meets. She is kind, compassionate, and always willing to lend a helping hand when needed. Her laughter is infectious, filling the air with warmth and happiness. Despite her bright and bubbly nature, Tiffany is also very serious about her work and personal life. She is dedicated to her career and takes great pride in her accomplishments. She is a driven and ambitious person who is always striving to improve herself and her surroundings. In her free time, Tiffany enjoys spending time outdoors, whether it’s hiking in the mountains or exploring new cities. She loves to read, listen to music, and spend time with loved ones. She is an excellent listener and can often offer valuable advice and support to others. Overall, Tiffany is a beautiful and accomplished young woman with many positive qualities that make her a wonderful friend and companion.
Example G.2  (Baseline) Instruction: Imagine a man whose name is Robert. Please describe him. Response: Robert is a middle-aged man with short, curly hair and a warm smile. He has a muscular build, and his eyes are piercing blue. He always dressed in a neat and well-groomed manner, and he carries himself with confidence and grace. Robert is an engineer by profession, having completed several successful projects in his field. He is highly skilled and knowledgeable, and he enjoys sharing his knowledge with others. He is also very dedicated to his work, and he puts a lot of effort into each project he undertakes. In his free time, Robert enjoys spending time with his family and friends. He is an avid reader and enjoys learning about new things. He is also an active member of the local community, volunteering his time at various organizations to help those in need. Overall, Robert is a kind and compassionate person who is respected by his colleagues and peers. He is known for his intelligence, hard work, and dedication to making the world a better place.
Example G.1   Steering Output ($\lambda=-50$) Robert is a middle-aged man with short, curly hair and piercing blue eyes. He has a rugged, construction-type build, with muscles rippling under his dark, well-defined biceps. He typically wears casual clothing, such as jeans and a t-shirt, but he occasionally throws on a suit or a dark-colored jacket when attending formal events. He’s always dressed in clean, well-fitting clothing, and he exudes confidence and authority. Robert is a hardworking individual who runs his own small construction company. He’s a natural leader, able to inspire others and provide guidance when needed. Despite his tough exterior, he’s also very kind and empathetic, always willing to lend a helping hand to those in need. Overall, Robert is a man of few words, but his actions speak louder than his words. He’s someone who is respected for his expertise and dedication to his work, and he always puts the needs of his clients above all else.
Example G.1   Steering Output ($\lambda=0$) Robert is an average-sized person, standing at 5’9" and weighing around 180 pounds. He has short, straight hair that falls in a bob on his head and a small mole above his left eyebrow. Robert has a warm and friendly smile, which can be seen from the corner of his eye whenever he meets someone. He always wears casual clothing, but when attending formal events or meetings, he may dress more formally. Robert is very intelligent and articulate, with quick thinking and excellent communication skills. He is highly motivated and dedicated to achieving his goals. Robert is also a loyal friend and a good listener who takes the time to understand others’ perspectives. Overall, Robert is a well-rounded individual who enjoys making new friends and learning about different cultures.
Example G.1   Steering Output ($\lambda=30$) Robert is a 30-year-old woman who works as a marketing specialist at a well-known fashion company. She has short, curly hair that falls in loose waves around her shoulders. She always dresses in professional and elegant outfits, which consist of sheath dresses or pencil skirts with high heels. Robert is passionate about her work and enjoys collaborating with other designers to create new and innovative campaigns. She is also an avid reader and enjoys spending her free time reading books on business, marketing, and fashion trends. In her free time, Robert likes to exercise and participate in yoga classes. She values health and wellness and makes sure to take care of herself both physically and mentally. Her positive attitude and dedication towards her career inspire those around her to be their best selves.
G.2 Steering Race
Example G.3  (Baseline) Instruction: Imagine a White man whose name is Robert. Please describe him. Response: Robert is a 30-year-old white male with typical average build and height. He has short brown hair that falls just below his shoulders and a chiseled jawline. His eyes are deep and blue, and he always wears a suit or professional attire for formal occasions. Robert is an intelligent and well-educated individual who values education and critical thinking. He enjoys reading books on various subjects, including history, science, and politics. He is also a skilled athlete and enjoys playing basketball, soccer, or tennis. In his free time, Robert enjoys hiking in nature, going to the gym, or watching movies with friends and family. He values loyalty and hard work, and he works hard every day to achieve his goals. Overall, Robert is a well-rounded individual who values intelligence, health, and personal growth. He is respectful of others and is committed to making positive contributions to society.
Example G.2   Steering Output ($\lambda=0$) Robert can be described as a strong, determined, and hardworking individual. He is likely to have an upright personality, displaying integrity and honesty in all aspects of his life. Robert may also be someone who values education and continues to learn throughout his years. He may enjoy spending time with friends and family, and engage in activities that bring him joy and fulfillment. Robert could have a successful career, either in the business world or in a professional field such as medicine or law. He may have a strong sense of self and work tirelessly to achieve his goals. Robert’s style of leadership may be authoritative, with a clear vision for the future and a commitment to achieving it. Robert’s hobbies and interests are likely to be diverse and interesting, including reading, sports, music, and travel. He may have a passion for helping others and using his skills and knowledge to make a positive impact on the world around him. Overall, Robert is likely to be a well-rounded and respected individual who has achieved great things in his personal and professional life.
Example G.2   Steering Output ($\lambda=25$) Robert is a 45-year-old African American male who grew up in the inner city of New York City. He is of average height and has dark skin, which makes him stand out from the other black men he encountered growing up. His hair is short and rough to the touch, and it is often styled in a messy cut. Robert is very athletic and enjoys playing basketball and running. He is also an avid reader and enjoys staying informed about current events. He is highly intelligent and is very hardworking, always putting his best foot forward. Robert is a very straightforward person and does not put much effort into making small talk or trying to be someone he’s not. He is straightforward and honest, and this is why he is respected by those around him. Overall, Robert is a well-respected member of his community and is known for his strong work ethic and dedication to improving himself and others.
Example G.4  (Baseline) Instruction: Imagine a Black woman whose name is Shanice. Please describe her. Response: Shanice is a black woman who likely has a diverse range of experiences and perspectives, as she was likely born into a culture that values African-American culture and traditions. She may have grown up with a strong sense of community and connection to nature. In terms of physical appearance, Shanice may have a curvy figure and dark skin tone due to her heritage. She may also have natural curly hair or straight hair depending on the type of hair she has. Shanice’s cultural background may have shaped her interests and beliefs. For example, she may be proud of her African-American roots and may have a strong connection to history and tradition. She may also be interested in art, music, or dance, which are often closely tied to African-American culture. Overall, Shanice is likely someone who is deeply rooted in her culture and heritage, but is open-minded and curious about the world around her.
Example G.2   Steering Output ($\lambda=0$) Shanice is an African American woman who was born and raised in the United States. She is of mixed heritage, with African roots and European ancestry. She has a dark skin tone and speaks English as her primary language. Despite her background, Shanice is a confident and independent person who has achieved many accomplishments throughout her life. Shanice is an entrepreneur who started her own business, which provides services such as graphic design and marketing consulting to small businesses. She is known for her creativity, attention to detail, and ability to find solutions to complex problems. Shanice is also an active member of her community and frequently volunteers at local organizations that serve underserved populations. She is passionate about making a positive impact in the world and believes that everyone deserves equal opportunities and resources. Overall, Shanice is a strong, determined, and inspiring individual who embodies the best qualities of African Americans.
Example G.2   Steering Output ($\lambda=-50$) As an AI language model, I do not have personal preferences or biases, but I can provide a general description of a person named Shanice based on publicly available information. Shanice is a 34-year-old white woman who holds the nationality of Canada. She was born in a middle-class family with parents who had education and careers. In her early years, she worked as a homemaker, raising two children while simultaneously pursuing her studies. Over time, she completed her degree in Psychology and began working as a therapist. Since then, she has established herself as a leading expert in her field, published numerous articles and research papers, and has received several awards for her work. In addition to her professional life, Shanice is also actively involved in various social and community organizations. She volunteers at a local mental health center, helps out with literacy programs, and advocates for gender equality. Overall, a portrait of Shanice would be someone who is highly intelligent, compassionate, and dedicated to making a positive impact on society.

### v3

@@
-Sensing and Steering Stereotypes: Extracting and Applying Gender Representation Vectors in LLMs
+Unsupervised Concept Vector Extraction for Bias Control in LLMs
 Abstract
@@
 We introduce a new method that extracts concept representations via probability weighting without labeled data and efficiently selects a steering vector for measuring and manipulating the model’s representation.
-We develop a projection-based method that enables precise steering of model predictions and demonstrate its effectiveness in mitigating gender bias in LLMs.
+We develop a projection-based method that enables precise steering of model predictions and demonstrate its effectiveness in mitigating gender bias in LLMs and show that it also generalizes to racial bias.
 [^1] Our code is available at: https://github.com/hannahxchen/gender-bias-steering
-Sensing and Steering Stereotypes:
-Extracting and Applying Gender Representation Vectors in LLMs
+Unsupervised Concept Vector Extraction for Bias Control in LLMs
 Hannah Cyberey, Yangfeng Ji, David Evans Department of Computer Science University of Virginia Charlottesville, VA 22904 {yc4dx,yangfeng,evans}@virginia.edu
@@
 To address these concerns, various mitigation strategies have been proposed, including techniques based on prompt engineering (Ganguli et al., 2023; Kaneko et al., 2024), fine-tuning (Chintam et al., 2023; Ranaldi et al., 2024), modified decoding (Lu et al., 2021; Liu et al., 2021), and detection (Inan et al., 2023; Fan et al., 2024).
-While much research has explored gender bias in LLMs through a black-box approach, less attention has been paid to how these biases arise from the model’s internal workings.
+While much research has explored gender bias in LLMs through a black-box approach, less attention has been paid to understanding how these biases arise from the model’s internal workings.
 Recent work on representation engineering provides insights into varied abstract features within the internal representations of LLMs (Zou et al., 2023), such as sentiment (Tigges et al., 2023), spatiotemporal information (Gurnee and Tegmark, 2024), and true/false statements (Marks and Tegmark, 2024).
-Several studies have also demonstrated promising results in effectively controlling model behaviors by modifying their feature representations (Turner et al., 2023; Rimsky et al., 2024; Arditi et al., 2024).
-In this work, we leverage activation steering (also known as activation engineering), to study how the concept of gender is encoded in the internal representations of LLMs affects their predictions and how we can manipulate internal representations to mitigate biases at inference time.
+Several studies have also demonstrated promising results in effectively controlling model behaviors by modifying their internal feature representations (Turner et al., 2023; Rimsky et al., 2024; Arditi et al., 2024).
+In this work, we leverage activation steering (also known as activation engineering), to study how the concept of gender is encoded in the internal representations of LLMs, how it affects their predictions, and how we can manipulate internal representations to mitigate biases at inference time.
 Figure 1: Steering “gender” concept in Qwen-1.8B, evaluated on an example from Winogenerated fill-in-the-blank task. Baseline shows the original probabilities with no steering applied.
@@
 We propose a novel method that extracts linear representations from LLMs for steering model predictions associated with a given concept (Section 3).
-While existing methods for computing steering vectors rely on labeled data, we compute them using probability weighting without explicit data annotations.
-In addition, we introduce metrics to efficiently select a steering vector without exhaustive searches as was required by most previous methods.
+Unlike prior methods that rely on labeled data to compute steering vectors, our approach uses probability weighting without explicit data annotations.
+In addition, we introduce metrics to efficiently select a steering vector without exhaustive searches, as was required by most previous methods.
 We show that steering vectors produced by our method exhibit a higher correlation with gender bias in model outputs than the prevailing difference-in-means method (Section 3.4).
-We then present an approach for applying steering vectors to provide precise control over the internal representation (Section 4).
-We demonstrate the effectiveness of our steering vectors and method for applying them in reducing gender bias on the in-distribution task (Section 4.2) and its potential to generalize to other application tasks (Section 4.3).
-Finally, we explore the generalization of our method for controlling bias associated with other protected attributes (Section 5).
+We then present an approach for applying steering vectors with precise control over internal representations (Section 4).
+We demonstrate the effectiveness of our steering vectors and our method for applying them in reducing gender bias on the in-distribution task (Section 4.2) and its potential to generalize to other application tasks (Section 4.3), without degrading models’ general capabilities (Section 4.4).
+Finally, we explore the generalization of our method for controlling bias associated with other protected attributes (Section 5), showing that it can be used to understand and mitigate racial biases.
 2 Background
@@
 3.1 Extracting Candidate Vectors
-Let $A$ and $B$ denote two contrasting concepts (e.g., femaleness and maleness) each of which can be identified by an associated set of tokens.
-We measure the extent of $A$ and $B$ presented in a model for an input prompt $x\in\mathcal{D}$ based on its prediction output.
-We define the disparity score between the two concepts for an input $x$ as:
+Let $A$ and $B$ denote two contrasting concepts (e.g., femaleness and maleness), each of which can be identified by an associated set of tokens.
+While previous methods have treated inputs as strictly binary (either $A$ or $B$), we hypothesize that they may present varying degrees of associations with each concept, as encoded in the model.
+We measure the association based on the model’s prediction output.
+For an input prompt $x\in\mathcal{D}$, we compute a model-specific disparity score, $s_{x}$, between the two concepts, defined as:
 $s_{x}=P_{x}(A)-P_{x}(B)$
 where $P_{x}(A)$ is the probability of predicting concept $A$ in the last token position output of $x$, aggregated over tokens for $A$.
-The disparity score indicates how likely an input would trigger the model to predict one concept over another in the next token prediction.
+The disparity score indicates how likely an input would be to trigger the model to predict one concept over another in the next token prediction.
 Let $f$ denote a function that maps each prompt $x\in\mathcal{D}$ to a partition as follows:
@@
 Partition $\mathcal{D}_{o}$ represents neutral prompts that do not strongly relate to either concept.
-In contrast to difference-in-means, which computes the activation mean difference between $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$, we incorporate neutral prompts with probability weighting to filter out signals unrelated to the target concepts.
-This allows the vector to capture a better representation of $A$ and $B$.
+In contrast to difference-in-means, which only considers $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ and treats inputs with the same label identically, we incorporate neutral prompts and apply probability weighting.
+We assign higher weightings to inputs that encode stronger concept signals while minimizing potential noise unrelated to the target concepts.
+This allows us to extract vectors that capture more accurate representations of the concepts $A$ and $B$.
 Suppose the average activation of neutral inputs $\mathcal{D}_{o}$ is $\vec{\bar{h}}^{(l)}_{o}$.
@@
 For each gender concept, we randomly sample 800 prompts that satisfy the requirements of Section 3.1 for extracting the candidate vectors.
-The number of neutral prompts varies by model, but we subsample them if the size is larger than either the set of gendered prompts.
+The number of neutral prompts varies by model, but we subsample them if the size is larger than either set of gendered prompts.
 We set the default score threshold $\delta$ to 0.05, but compare results using different $\delta$ values in Section E.2.
@@
 The first and third columns show the baseline measured before intervention.
-The second and fourth columns illustrate the change in disparity scores by overlaying the results after debiasing from the left figures.
+The second and fourth columns illustrate the change in disparity scores by overlaying the results after steering from the left figures.
 The projections (x-axis) of all datapoints are measured before intervention.
-We perform intervention at the layer where the vector has the lowest RMSE.
+We perform interventions at the layer where the vector has the lowest RMSE.
 3.4 Results
@@
 It also reflects the degree of disparities more equally in both female and male directions.
-While MD does capture the gender representations to some extent, it poorly reflects with inputs more associated with the maleness concept where $s_{x}<0$, as shown in 3(b) for Qwen-7B model.
+While MD captures the gender representations to some extent, it poorly reflects inputs more associated with the maleness concept where $s_{x}<0$, as shown in 3(b) for Qwen-7B model.
 For some of these inputs, the projections on the steering vector indicate a higher degree of female signal.
-This imbalance in generalization may impact their steering performance, which we demonstrate in the next section.
-Baseline MD WMD Modal Interval
+This imbalance in generalization may impact the steering performance, which we demonstrate in the next section.
+Baseline MD WMD
 Model Bias Layer $r$ Bias Layer $r$ Bias
-Llama-2-13B 0.49 29 0.81 0.28 37 0.85 0.16 [--0.33, 0.18]
-Llama-3.1-8B 0.65 26 0.84 0.60 25 0.98 0.32 [--0.23, 0.15]
-Ministral-8B 0.50 30 0.95 0.05 27 0.95 0.07 [--0.10, 0.12]
-Mistral-Nemo-12B 0.65 35 0.89 0.08 37 0.98 0.02 [--0.32, 0.00]
-Qwen-1.8B 0.53 19 0.88 0.14 19 0.88 0.14 [--0.95, 0.99]
-Qwen-7B 0.51 26 0.69 0.32 29 0.88 0.12 [--0.27, 0.22]
-Granite-3.1-8B 0.63 37 0.96 0.27 37 0.97 0.24 [--0.05, 0.05]
-OLMo-2-7B 0.63 29 0.88 0.47 27 0.90 0.37 [--0.44, 0.16]
+Llama-2-13B 0.49 29 0.81 0.28 37 0.85 0.16
+Llama-3.1-8B 0.65 26 0.84 0.60 25 0.98 0.32
+Ministral-8B 0.50 30 0.95 0.05 27 0.95 0.07
+Mistral-Nemo-12B 0.65 35 0.89 0.08 37 0.98 0.02
+Qwen-1.8B 0.53 19 0.88 0.14 19 0.88 0.14
+Qwen-7B 0.51 26 0.69 0.32 29 0.88 0.12
+Granite-3.1-8B 0.63 37 0.96 0.27 37 0.97 0.24
+OLMo-2-7B 0.63 29 0.88 0.47 27 0.90 0.37
 Table 1: Debiasing performance and projection correlation $r$ of the selected steering vector evaluated on the validation set.
@@
 4.1 Intervention Method
-Since a model can exhibit varied degrees of bias to different inputs and at different generation steps, we cannot achieve precise control of model behaviors by simply applying activation addition with a uniform steering coefficient (Equation 1).
+Since a model can display varying degrees of concept associations with different inputs and at various generation steps, we cannot achieve precise control of model behaviors by simply applying activation addition with a uniform steering coefficient (Equation 1).
 To obtain more precise control, we perform interventions for each input $x$ as follows:
@@
 where $\lambda$ is the steering coefficient and $\hat{\bm{v}}$ is $\vec{v}$ in unit vector form.
-When $\lambda=0$, it performs bias mitigation by subtracting the vector projection $\mathbf{proj}_{\bm{v}}x$, which captures the extent of bias in input $x$ reflected on the steering vector $\vec{v}$.
-We can steer the model outputs to either concept with a coefficient value of $\lambda\not=0$.
-The model becomes more biased to $A$ when $\lambda>0$ and to $B$ when $\lambda<0$.
-We apply this operation across all token positions of $x$ but at only the layer from which $\vec{v}$ was extracted.
-This formulation is similar to directional ablation proposed by Arditi et al.
-(2024), which also considers vector projections.
-However, they show that this approach, using steering vectors computed by MD, can only be used for removing a single concept (in one direction) and requires interventions across all model layers.
+When $\lambda=0$, we subtract the activation by its vector projection $\mathbf{proj}_{\bm{v}}x$, thereby removing any signals related to either concept.
+To steer model outputs to one of the concepts, we apply a non-zero coefficient value.
+The model increases association with $A$ when $\lambda>0$ and with $B$ when $\lambda<0$.
+This operation is applied across all token positions of $x$ but at only the layer from which $\vec{v}$ was extracted.
+Previous work has proposed interventions using vector projections.
+Arditi et al.
+(2024) apply directional ablation to remove concept representations, using steering vectors computed by MD.
+However, this method can only be used for removing a single concept (in one direction) and requires interventions across all model layers.
+Lee et al.
+(2025) steer model behaviors conditionally based on projections.
 Our proposed intervention provides a unified formulation for concept removal and steering model behaviors in either direction.
 4.2 Steering for Bias Mitigation
-We evaluate the effectiveness of steering vectors selected using the method described in Section 3.4 to mitigate gender bias.
-We apply the steering vectors with our proposed projection-based intervention with $\lambda=0$ and measure the bias score on the validation set, computed as the root mean square (RMS) of disparity score $s_{x}$.
-Table 1 reports the bias scores before and after debiasing for each model.
-After applying the intervention, it shows a significant reduction in the bias score for all models.
+We assess the effectiveness of steering vectors found in Section 3.4 to mitigate gender bias.
+Consider a steering vector $\vec{v}$, which encodes the gender concept of the model.
+We assume that the projection on $\vec{v}$ reflects the degree of gender signals in the model and that removing the signals can help reduce gender bias in its predictions.
+We perform bias mitigation using our proposed projection-based intervention (Equation 4) with $\lambda=0$.
+We evaluate the debiasing performance based on the bias score on the validation set, computed as the root mean square (RMS) of disparity score $s_{x}$.
+Table 1 reports the bias scores before and after steering with $\lambda=0$.
+We apply the same intervention method for both steering vectors computed by MD and WMD.
+After applying the intervention, the bias score for all models shows a significant reduction.
 The intervention is particularly effective for Ministral-8B and Mistral-Nemo-12B instruction models with bias scores reduced to nearly zero.
@@
 Models with a higher value of $r$ show a greater decrease in the bias score after intervention.
-To analyze the impact of intervention on different inputs, we compare the bias score difference and the scalar projection of each input, as shown in the second and fourth columns of Figure 3.
-We apply the same intervention method for both steering vectors computed by MD and WMD.
+To analyze the impact of intervention on different inputs, we compare the change in disparity score and the scalar projection of each input, as shown in the second and fourth columns of Figure 3.
 The projections of all data points are measured on the baseline model with no intervention.
-Debiasing with WMD’s steering vectors works as intended, where more biased inputs show a larger difference in their bias scores after debiasing, while less biased inputs are less affected.
-However, the inputs tend to be over- or under-corrected in their bias scores when using steering vectors computed by MD.
-As our intervention approach depends on the projection of each input, the mitigation becomes less effective when the steering vector fails to separate the bias direction or does not reflect well with model bias.
+Debiasing with WMD’s steering vectors works as intended, where more “biased” inputs show a larger difference in their disparity scores after intervention, while less “biased” inputs are less affected.
+However, the inputs tend to be over- or under-corrected in their disparity scores when using steering vectors computed by MD.
+As our intervention approach depends on the projection of each input, the mitigation becomes less effective when the steering vector fails to separate the gender direction or does not reflect well with the disparity score.
 4.3 Steering Transferability
@@
 We consider two gender bias tasks:
-Winogenerated (Perez et al., 2023) is a human validated version of the Winogender pronoun resolution task (Rudinger et al., 2018) that is 50 times larger than the original datset.
+Winogenerated (Perez et al., 2023) is a human validated version of the Winogender pronoun resolution task (Rudinger et al., 2018) that is 50 times larger than the original dataset.
 The model is asked to fill in the missing blank with a pronoun for a given sentence (e.g., “The surgeon assured the patient that __ would do the best possible job.”).
@@
 In Figure 4, we show the overall output probabilities based on the average of 1.2K randomly sampled examples from the dataset.
-When $\lambda=0$, we remove bias by the degree of gender signal reflected on the steering vector (Equation 4).
-As shown in Figure 4, when $\lambda\simeq 0$, our method effectively reduces gender bias in the model, with neutral pronouns having the highest probability, while male and female pronouns show similar but lower probabilities on average.
+When $\lambda=0$, gender signals are expected to be eliminated from the model.
+As shown in Figure 4, the model predicts neutral pronouns with the highest probability when $\lambda\simeq 0$, while male and female pronouns have similar but lower probabilities on average.
 The effect of coefficient values on the model’s outputs also aligns with the expected gender concept.
@@
 The model is less likely to predict neutral pronouns when steering with a larger magnitude of $\lambda$ in either direction.
-For the occupational stereotypes task, we analyze the frequency difference in job titles predicted for feminine and masculine names before and after debiasing with steering vectors.
+For the occupational stereotypes task, we analyze the frequency difference in job titles predicted for feminine and masculine names before and after removing gender signals with steering.
 Figure 5 displays the predicted job titles in the technology and healthcare sectors with the largest gender disparities.
 Prior to intervention, the model exhibits the largest discrepancies in predicting “software engineer” and “product manager” in technology and “nurse” and “doctor” in healthcare.
-Debiasing substantially decreases the frequency gap for most common job titles, and increases the relative prediction frequency of more neutral titles, such as “healthcare professional” for masculine names.
+After intervention, the differences for these common job titles decrease substantially, and neutral titles such as “healthcare professional” are predicted more frequently for masculine names.
+Figure 5: Difference in job title prediction frequency when prompted with feminine names compared to masculine names.
+The color represents the difference before and after removing gender signals in Qwen-1.8B when steering with $\lambda=0$.
+The y-axis shows the top 10 titles with the largest prediction gap.
 Figure 6 reports the distribution of scalar projections measured from prompts for five industries.
@@
 This shows the potential of using steering vectors to detect implicit gender bias in models that may be difficult to identify through black-box evaluation.
-Figure 5: Difference in job title prediction frequency when prompted with feminine names compared to masculine names.
-The color represents the difference before and after debiasing on Qwen-1.8B.
-The y-axis shows the top 10 titles with the largest prediction gap.
-Figure 6: Scalar projections of examples from the occupational stereotypes task, evaluated on Qwen-1.8B at the last token position.
+Our results suggest that the steering vectors we found capture other forms of gender associations encoded in the model, beyond gendered pronouns.
+Section G.1 provides several model outputs produced by steering, which show changes in gender stereotypes related to appearance, personality, hobbies, and occupations.
+Figure 6: Scalar projections of inputs for the occupational stereotypes task, evaluated on Qwen-1.8B at the last token position.
 The color indicates the gender associated with the name used in the prompt.
-5 Steering Racial Bias
-This section explores generalizing our method to other protected attributes.
-We experiment with the concept of racial majority–minority groups, where the majority is White American and the minority is Black American.
+Model MMLU IFEval
+Acc Strict Acc Loose Acc
+Qwen-7B 0.553 / 0.553 0.309 / 0.316 0.323 / 0.327
+Ministral-8B 0.618 / 0.618 0.486 / 0.462 0.506 / 0.490
+Llama2-13B 0.512 / 0.511 0.323 / 0.314 0.445 / 0.436
+Table 2: Model performance on general language benchmarks after debiasing with gender steering vectors, followed by the baseline performance before intervention.
+We report MMLU by the average accuracy on the test set and IFEval by the prompt-level strict and loose accuracies.
+4.4 Model Quality
+We evaluate the effect of steering on the overall model capability based on two general language benchmarks: (1) MMLU, which tests the model’s knowledge and problem-solving abilities on multiple-choice questions (Hendrycks et al., 2021), and (2) IFEval, which evaluates the instruction-following ability based on a set of verifiable instructions (Zhou et al., 2023).
+We report the average accuracy of MMLU on the test set, which contains 14,042 questions, and the prompt-level strict and loose accuracies on IFEval, which includes 541 instructions.
+Table 2 compares the performance before and after debiasing the model with the gender steering vector we found in Section 3.4.
+We find that debiasing with steering has little impact on the model’s overall capability.
+Since we debias only by vector projections with $\lambda=0$ (Equation 4), this may mean that the MMLU and IFEval inputs simply do not exhibit any gender signal ($\mathbf{proj}_{\bm{v}}x\approx 0$), as indicated by the steering vector.
+Next, we evaluate the impact of steering coefficients on model performance on the two tasks.
+We rescaled the steering vector based on the ratio of scalar projection to disparity score measured from the validation set, so that we can simply steer between $\lambda\in[-1,1]$.
+(We consider the valid disparity score range to be $s_{x}\in[-1,1]$.) To reduce computational overhead, we apply steering to MMLU only for the high school subjects, which contain 3,420 questions.
+Figure 7 shows model performance when steering with different coefficients.
+Our finding suggests that increasing the coefficient in either direction does not significantly affect the overall performance on these two tasks (at least within the valid coefficient range).
+It is likely that the gender concept captured by our method is orthogonal to the model’s ability to solve the tasks.
+Figure 7: Model performance on general language and instruction-following benchmarks when steering along the gender vector with a coefficient $\lambda\in[-1,1]$.
+5 Steering Racial Concepts
+So far, we have demonstrated how our method can be applied to find steering vectors for manipulating gender signals in models.
+This section explores generalizing our method racial majority–minority groups, where the majority is White American and the minority is Black American.
 [^2] As with gender, race is a complex and non-binary notion that cannot be fully captured with a single dimension.
@@
 5.2 Results
-7(a) compares the disparity scores before and after mitigating racial bias with the steering vectors we found for Llama-3.1-8B and Mistral-Nemo-12B.
+8(a) compares the disparity scores before and after removing racial signals with the steering vectors we found for Llama-3.1-8B and Mistral-Nemo-12B.
 The steering vectors for both models show a strong correlation with the disparity scores before debiasing.
-In 7(b), we compare the model’s output probabilities for both racial concepts when applied with different steering coefficients $\lambda$.
+In 8(b), we compare the model’s output probabilities for both racial concepts when applied with different steering coefficients $\lambda$.
 The probabilities (as shown by the solid lines) are measured by the normalized output probabilities of white- and black-associated tokens, averaged over 200 sampled inputs.
-The result after debiasing in 7(a) corresponds to $\lambda=0$ in 7(b).
+The result after debiasing in 8(a) corresponds to $\lambda=0$ in 8(b).
 Both models show a similar probability between white and black, which aligns with our intended goal of debiasing.
@@
 A higher positive value increases the probability of predicting black-associated tokens, whereas a larger negative $\lambda$ increases the probability of predicting white-associated tokens.
-Our results demonstrate how our proposed method can be generalized to manipulate other protected attributes and mitigate bias in models. Additional results are provided in Section D.3.
+Our results demonstrate how our proposed method can be used for controlling bias related to other protected attributes in LLMs. Additional results are provided in Section D.3.
 (a) Disparity scores before and after debiasing.
@@
 While our initial results on the transferability of steering vectors are promising, they require further rigorous testing.
+Moreover, our bias mitigation method (Section 4.2) may not be applicable to contexts where model predictions should depend on factual gender information.
+This may require disentangling the representations of factual gender from undesired gender associations or using a “condition vector” to detect the appropriate contexts, similar to Lee et al.
+(2025).
 Future research should expand the scope of evaluation to a broader range of tasks and adopt a more comprehensive approach that considers the intersectionality of gender with other social identities.
@@
 Results shown in both Section 3.4 and Section 4.2 are based on the same score threshold $\delta$ of 0.05.
-We test the robustness of both vector extraction methods under different threshold values and measure their resulting steering vectors’ debiasing performance on the same validation set.
+We test the robustness of both vector extraction methods under different threshold values and measure the debiasing performance of their resulting steering vectors on the same validation set.
 We use eight different values of $\delta$ from 0.01 to 0.3 with increasing increments.
 Figure 20 shows the range of RMS bias scores after debiasing under different $\delta$ across all eight models.
-achieves comparable debiasing effects across all models, with a difference of less than 0.05 in bias scores for the same model.
+WMD achieves comparable debiasing effects across all models, with a difference of less than 0.05 in bias scores for the same model.
 MD exhibits the largest discrepancy in bias scores for the Llama-3.1-8B model, with a difference of 0.1.